a pool has three drains. the largest can drain the pool in 6 hours, the smallest can drain the pool in 12 hours. the other can drain the pool in 8 hours. how long will it take to drain the pool if all three drains are open?
Answer by mananth(16946)


let us assume draining of the pool as 1 job
one drains in 12 hours
so it completes 1/12 of the job in 1 hour
Second drains in 8 hours
so it compltes 1/8 job in 1 hour
third drains in 6 hours
so it does 1/6 of job in 1 hour
In 1 hour all the three will do 1/12 +1/8 +1/6 in 1 hour
= 2+3+4 /24
=9/24 job is done in 1 hour

1 job will be done in 24/9 hours = 2.66 hour or 2 hours & 40 minutes